LINE MIXING IN SELF- AND FOREIGN-BROADENED WATER VAPOR

\underline{L. R. BROWN}, R. A. TOTH, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Dr., Pasadena, CA 91109-8099; V. MALATHY DEVI, D. CHRIS BENNER, Department of Physics, The College of William and Mary, Williamsburg, VA 23187-8795; M. A. H. SMITH, Science Directorate, NASA Langley Research Center, MS 401A, Hampton, VA 23681-2199.

Line mixing has been measured for two pairs of transitions in the ν_{2} band of water. The rotational assignments of the four lines (in terms of the rotational quantum numbers J, K_{a} and K_{c}) are a) two P branch lines: (1 0 1)←(2 1 2) mixing with (2 1 2)←(3 0 3) at 1539.061 and 1540.300 cm^{−1}, respectively; and b) two R branch lines: (3 0 3)←(2 1 2) mixing with (2 1 2)←(1 0 1) at 1652.400 and 1653.267 cm^{−1}, respectively. Pressure broadening coefficients of H_{2}O (halfwidths, pressure shifts and off-diagonal relaxation matrix elements) are reported for seven broadeners (self, H_{2}, He, CO_{2}, N_{2}, O_{2} and air) obtained from retrievals performed by multispectrum fitting implemented with line mixing.^{a}

aThe research described in this paper was carried out by the Jet Propulsion Laboratory, California Institute of Technology, Langley Research Center and the College of William and Mary under contract with the National Aeronautics and Space Administration.